OpenStudy (anonymous):

HELP write a polynomial function with the zeros: ( write in standard form ) 4,-8, and 2+5i

OpenStudy (anonymous):

is that 4-8

OpenStudy (precal):

(x-4)(x+8)(x-(2+5i))(x-(2-5i))

OpenStudy (precal):

just multiply it out

OpenStudy (anonymous):

I tried that but the answer I got didn't equal any of the possible answer choices; is there any other way that I could solve the problem

OpenStudy (anonymous):

is it 4-8 or 4,-8

OpenStudy (precal):

I am sure that is 4 and -8 are two of the zeros

OpenStudy (anonymous):

4 is a zero -8 is a zero and 2+5i is a zero

OpenStudy (precal):

don't forget to assume 2-5i as a zero

OpenStudy (anonymous):

my possible solutions are: A. f(x)= x^4- 6x^3- 20x^2+ 122x- 928 B. f(x)= x^4 - 19x^2 +244x -928 C. f(x)=x^4- 6x^3+ 20x^2- 122x+928 D. f(x)= x^4 -61x^2+ 244x -928

OpenStudy (anonymous):

yeah I remembered that, this is for a home work assignment on an online class, not a quiz, but I really need to understand it. I understood what @precal said by multiplying the zeros but I haven't been able to correctly match my answer to any of the choices

OpenStudy (precal):

http://www.mathwarehouse.com/algebra/complex-number/multiply-complex-number.php here this might help gotta go